What is the distance formula?
How do you solve the problem: x-3y=17
2x+y=6 (using the sustitution method)
A cable 50ft long is cut into two pieces so that one piece is 6ft longer than the othe. How long is each piece?

Answer by checkley71(8403)

THE DISTANCE FORMULA FOR VEHICLES/BOATS/PLANES/BICYCLES/ETC. IS
DISTANCE = RATE*TIME
EX. MILES=45 MPH*3 HOURS
MILES=45*3
MILES=135
-----------------------------------------------------------------
X-3Y=17 OR X=3Y+17
2X+Y=6 SINCE X=(3Y+17) WE SUBSTITUTE (3Y+17) FOR X IN THIS FORMULA
2(3Y+17)+Y=6
6Y+34+Y=6
7Y=6-34
7Y=-28
Y=-28/7
Y=-4 SOLUTION. NOW REPLACE Y WITH -4 AND SOLVE FOR X
X-3*-4=17
X+12=17
X=17-12
X=5 SOLUTION
PROOF
2*5-4=6
10-4=6
6=6